dakotagallen:

Octavian became the emperor of Rome in 27 B.C. He ruled until his death in A.D. 14. How long did he rule as emperor? (Hint: Remember there was no A.D. 0.) 1.40 years 2.13 years 3.41 years 4.12 years

jhonyy9:

1 year 10 m x year 410 m x = 410/10 = 41 41/12 = 3,41 years

jhonyy9:

in this time 1 year was considered that has just 10 month

jhonyy9:

and the question is in year with 12 month

jhonyy9:

hope now this is more understandably sure yes ?